Find two consecutive odd integers such that 50 more than the lesser is five times the greater.
Found 2 solutions by ikleyn, Edwin McCravy:
Answer by ikleyn(53487)

.
Find two consecutive odd integers such that 50 more than the lesser
is five times the greater.
~~~~~~~~~~~~~~~~~~~~~~~~

Let the numbers be x and (x+2).

Then we can write this equation

    x + 50 = 5*(x+2).


Simplify and find x

    x + 50 = 5x + 10

    50 - 10 = 5x - x

       40   =   4x

        x   =   40/4 = 10.


It shows that the problem is posed INCORRECTLY.

Such odd integers as described in your post DO NOT EXIST,
and the problem describes a situation which never may happen.

Solved.

======================

It is a grave sin to spread wrong tasks in the Internet.



Answer by Edwin McCravy(20077)

Change 'even' to 'odd' and you can do it.

Find two consecutive EVEN integers such that 50 more than the lesser
is five times the greater.

   G = L+2
L+50 = 5G

L+50 = 5(L+2)
L+50 = 5L+10
  40 = 4L
  10 = L
   G = L+2
   G = 10+2
   G = 12

10 and 12.

Or switch "greater" and "lesser" and you can do it.

Find two consecutive odd integers such that 50 more than the GREATER
is five times the LESSER.

     G = L+2
  G+50 = 5L

L+2+50 = 5L
  L+52 = 5L
    52 = 4L
    13 = L
     G = L+2
     G = 13+2
     G = 15.

13 and 15.

There is no solution as you have written it.  Maybe your teacher
was testing you to see if you'd know it has no solution.  Sometimes
the correct answer is "no solution".

Edwin